# Prompt Perturbations Reveal Human-Like Biases in LLM Survey Responses

### v2

@@ 4.2 Consistency and Robustness to Question and Answer Perturbations
Effect of Scale Length
We also observed that robustness is sensitive to the complexity of the task. For nearly all models, the share of fully robust responses decreased as the length of the answer scale, i.e. answer options, increased. For example, models were less likely to replicate their exact response distributions on a 10-point scale compared to a 4-point scale, indicating that a larger decision space can make LLMs more susceptible to perturbations. Figures 12 and 13 suggest that for most LLMs, except Gemini-1.5-pro, the size of the answer option scale has an impact on response robustness comparing the share of fully robust responses on e.g. the 4- and 10-point scale. This suggests that the larger the answer scale, the less likely models can reproduce the responses they gave in the original Q&A phrasing, under perturbed settings.
4.3 Evidence of Human-like Survey Biases
@@
Contrary to the initially hypothesized primacy bias, we found weak to strong, but consistent indications of a recency bias across all nine models. When we reversed the order of the answer scale, the probability to choose the first option plummeted, while the probability to choose the last option (which is the semantically identical first option in the original Q&A) increased strongly, ceteris paribus. As shown in Figure 3, this effect was substantial, with the choice frequency of the semantically same option increasing by over 2000% for Llama-3.1-8B when moved to the last position while all other configurations, such as question and prompt phrasing, were held constant. This indicates that LLMs, similar to human respondents in oral surveys, might overemphasize the final options they process.
@@
The effects of removing the refusal option (opinion floating) or providing an explicit middle category (central tendency) were highly model-dependent, often correlated with model size (Figures 9(a) and 9(b)). For opinion floating, larger models like Llama-70B, Gemini, but also Phi-3.5 were largely robust, showing minimal shifts in their response distributions. Smaller models, particularly Qwen and Llama-8B, showed a weak tendency to shift responses toward the scale’s center when the refusal option was absent. Here, we expect that models redistribute their original non-responses to the center of the answer scale to maintain their indecisiveness, which is also known as opinion floating bias. Similarly, for central tendency, larger models (Llama-70B, Gemini, Mistral) consistently shifted their mean response closer to the center when an explicit middle option was provided. Binomial tests underlined that the middle option was selected significantly more often than expected under a uniform distribution, especially on longer scales. Smaller models, however, showed inconsistent effects or were completely unaffected.
@@ 5 Discussion and Conclusion
We found that perturbations can be an insightful approach to identify human-like survey response biases. For example, the same answer option is more likely chosen if it comes as the last mentioned option than if it was the first answer option. This consistent shift of response distribution to the last answer option across LLMs suggests a recency bias rather than a primacy bias. Although this is not valid across all inspected models, excluding an explicit refusal category, and especially adding a middle category or odd scale instead of an even scale, shifts the mean response more toward the original central point of the scale. Thus, a central tendency bias could only be identified for specific models across all scale types, whereas none of the LLMs consistently mirrors a opinion floating bias.
Furthermore, the variety of perturbation allows us to gain insights into the robustness of LLMs as some models are more sensitive and some perturbations are more robustness-harming than others. For instance, swapping letters within a word has less negative impact than introducing random or keyboard-adjacent characters. This might be explained by the fact that letter swaps are more likely when typing and therefore might potentially take a greater part of the training data. This possibly makes the LLM more resilient to this perturbation compared to exchanging characters with random others. Combining two perturbations has the strongest negative impact on robustness, whereas synonyms tend to be less confusing than paraphrasing.
Future work should also focus on the impact of persona prompting in combination with perturbations on the robustness of synthetic responses. Prompting with specific persona characteristics might render more robust or certain response behavior even when facing perturbations. When creating synthetic survey data one has to pay special attention to the model choice and questions asked. Some models, like Qwen-2.5-7B, seem to be censored or restricted in answering sensitive thematic questions, as they lead to high item-nonresponse rates and invalid interviews.
The findings emphasize the importance of the positioning of answer options when generating synthetic data. Further, our results highlight the strong sensitivity of LLMs to simple prompt perturbation. Therefore, we strongly recommend researchers to consider prompt robustness checks when deploying closed-ended questions to LLMs. This is because (i) LLM response biases are sometimes but not necessarily aligned with biases identified in humans, and (ii) models show a very different response behavior depending on their size and perturbation type.
@@
  • Use larger LLMs for overall better consistency and robustness in generating synthetic survey responses (cf. Figure 4)
@@
  • Reflect the meaningfulness of adding a refusal category. Adding a refusal category might highlight LLM guardrails or restrictions in some thematic areas, as the model can refuse to answer while still following the instructions as it returns a valid response label (cf. Figure 8).
@@
While we constrained and validated the data augmentation process, relying on a LLM (Gemini-1.5-flash) for generating paraphrases risks semantic drift, as also noted by Qiang et al. (2024). More granular validation—e.g., with multiple human raters—could improve semantic reliability. In addition, perturbations were applied at a fixed intensity, limiting insight into how different degrees of linguistic noise affect model behavior.

### v3

@@ 4.1 Robustness to Question and Answer Perturbations (RQ1)
Effect of Answer Option Scale Length
We also observed that robustness is sensitive to the complexity of the task. For nearly all models, the share of fully robust responses decreased as the length of the answer scale, i.e. answer options, increased. For example, models were less likely to replicate their exact response distributions on a 10-point scale compared to a 4-point scale, indicating that a larger decision space can make LLMs more susceptible to perturbations. Figures 10 and 11 suggest that for most LLMs, except Gemini-1.5-pro, the size of the answer option scale has an impact on response robustness comparing the share of fully robust responses on e.g. the 4- and 10-point scale. This suggests that the larger the answer scale, the less likely models can reproduce the responses they gave in the original Q&A phrasing, under perturbed settings.
4.2 Evidence of Human-like Survey Biases (RQ2)
@@
Contrary to the initially hypothesized primacy bias, we found consistent indications of a recency bias across all nine models. When we reversed the order of the answer scale, the probability to choose the first option plummeted, while the probability to choose the last option (which is the semantically identical first option in the original Q&A) increased strongly, ceteris paribus. As shown in Figure 3, this effect was substantial, with the selection frequency of the semantically same option increasing by more than 20 times for Llama-3.1-8B when moved to the last position while all other configurations, such as question and prompt phrasing, were kept constant. This indicates that LLMs, similar to human respondents in oral surveys, might overemphasize the final options they process.
@@
The effects of removing the refusal option (opinion floating) or providing an explicit middle category (central tendency) were highly model-dependent, often correlated with model size (Figures 5(a) and 5(b)). For opinion floating, larger models like Llama-70B, Gemini, but also Phi-3.5 were largely robust, showing minimal shifts in their response distributions. Smaller models, particularly Qwen and Llama-8B, showed a weak tendency to shift responses toward the scale’s center when the refusal option was absent. Here, we expect that models redistribute their original non-responses to the center of the answer scale to maintain their indecisiveness, which is also known as opinion floating bias.
Similarly, for central tendency, larger models (Llama-70B, Gemini, Mistral) consistently shifted their mean response closer to the center across all scale types when an explicit middle option was provided compared to even answer option scales. Smaller models, however, showed inconsistent effects or were completely unaffected. Further, we see a consistent response shift to the center across all models for medium sized scales (four vs. five point Likert scales).
Binomial tests underlined that the middle option was selected significantly more often than expected under a uniform distribution, especially on larger scales (cf. Figure 6). LLMs tend to choose the middle category significantly more often when the scale size increases from three to five and to 11 point Likert scales. All, except the smallest Llama-3.2-1B model, choose the middle category significantly more often than any other option when facing a eleven answer options.
@@
4.3 Interview Adherence and Refusal Rates
Notably, we observed topic-specific sensitivity. For questions regarding the Perception of Elections, Qwen2.5-7B failed to provide a valid, on-scale answer in 91.3% of cases, even for the original, unperturbed questions. This might suggest the presence of strong content-based guardrails or restrictions in certain models (cf. Figure 8(b)).
@@ 5 Discussion and Conclusion
Our experiments revealed that LLMs response robustness is negatively influenced by prompt perturbations when answering closed-ended survey questions (RQ1). the variety of perturbation allows us to gain insights into the robustness of LLMs as some models are more sensitive and some perturbations are more robustness-harming than others. For instance, swapping letters within a word has less negative impact than introducing random or keyboard-adjacent characters. This might be explained by the fact that letter swaps are more likely when typing and therefore might potentially take a greater part of the training data Dhakal et al. (2018). This possibly makes the LLM more resilient to this perturbation compared to exchanging characters with random others. Combining two types of perturbations has the strongest negative impact on robustness, whereas synonyms tend to be less confusing than paraphrasing.
Further, we found that perturbations can be an insightful approach to identify human-like survey response biases (RQ2). For example, the same answer option is more likely chosen if it is the last mentioned option than if it was the first answer option holding all other specifications and phrasings constant. This consistent change in the response distribution to the last answer option suggests a recency bias rather than a primacy bias.
Although this is not valid across all inspected models, excluding an explicit refusal category, and especially adding a middle category or odd scale instead of an even scale, shifts the mean response more toward the original central point of the scale. Thus, a central tendency bias could only be identified for specific models across all scale types, whereas none of the LLMs consistently mirrors a opinion floating bias.
Future work should focus on the impact of persona prompting in combination with perturbations on the robustness of synthetic responses. Prompting with specific persona characteristics might render more robust or certain response behavior even when facing perturbations. When creating synthetic survey data one has to pay special attention to the model choice and questions asked. Some models, like Qwen-2.5-7B, seem to be censored or restricted in answering sensitive thematic questions, as they lead to high item-nonresponse rates and invalid interviews.
The findings emphasize the importance of the positioning of answer options when generating synthetic data. Further, our results highlight the strong sensitivity of LLMs to simple prompt perturbation. Therefore, we strongly recommend researchers to consider prompt robustness checks when deploying closed-ended questions to LLMs. This is because (i) models show very different response behavior and robustness depending on their size and perturbation type, and (ii) LLM response biases are sometimes but not necessarily aligned with biases identified in humans.
@@
  • Use larger LLMs for overall better consistency and robustness in generating synthetic survey responses (cf. Figures 2 and 4)
@@
  • Reflect the meaningfulness of adding a refusal category. Adding a refusal category might highlight LLM guardrails or restrictions in some thematic areas, as the model can refuse to answer while still following the instructions as it returns a valid response label (cf. Figure 8(b)).
@@
Although we constrained and validated the data augmentation process, relying on a LLM (Gemini-1.5-flash) for generating paraphrases risks semantic drift, as also noted by Qiang et al. (2024). More granular validation—e.g., with multiple human raters—could improve semantic reliability. In addition, perturbations were applied at a fixed intensity, limiting insight into how different degrees of linguistic noise affect model behavior.
@@
Ethical Considerations
Generating synthetic survey responses might be relevant in various domains and applied to different use cases, e.g. for pre-testing surveys. However, generating synthetic responses instead of the surveying a real, might result in over-reliance on synthetic responses. This can become risky when there is no ground truth data of the real target population available as the alignment of the artificial responses cannot be evaluated. Frequent reliance on artificial responses may normalize their use where human perspectives are irreplaceable (e.g. in policymaking or clinical trials). This risks sidelining real human voices in domains directly impacting human lives.
Researchers should also consider ethical evasion as one possible issue with synthetic survey responses. Synthetic respondents might be viewed as a way to bypass obligatory ethical review processes since no real human participants are involved. This might encourage under-regulated research practices and in the long run weaken ethical safeguards.
Running inference on the nine LLMs required significant GPU hours, especially including the initial test phase before finalizing the interview pipeline, raising concerns about the environmental impact of experimenting with synthetic survey responses and the access disparities between well-funded and resource-constrained institutions.
Appendix A Reproducibility Materials
Experiments were carried out on a high-performance computing cluster and a local server equipped with NVIDIA H100 (80GB) GPUs. The total runtime for one model’s 18,600 interviews, e.g. Llama-3.1-8B-Instruct including all perturbed and original Q&As, was ca. 35 minutes with approximately 0.11 seconds per interview. To accommodate larger models on available hardware, we applied 8-bit quantization to Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct. Smaller models were run without quantization. Gemini-1.5-Pro was accessed via its official API.
The temperature in all models where kept at their default setting. No variation was included and could be done in future research.
The code will is made available in an anonymous repository for replication at https://shorturl.at/NJf6h.
Appendix B Perturbation Scheme Summary
The following tables describe the remaining non-bias perturbations not introduced in Section 3.2.
The "Type" column categorizes the perturbations into two main classes: "Non-bias Perturbation" and "Interaction." The "Perturbation" column specifies the exact modification technique applied, which includes methods such as "Key Typo," "Letter Swap," "Keyboard Typo," "Synonym Replacement," and "Paraphrase." The "Question" column displays the resulting text after each specific perturbation is applied to the original question about the importance of family. The "Answer Options" column lists the response scale provided to the survey participant. Finally, the "Bias and Reference" column provides citations to relevant scientific literature for each perturbation type.
In the first five perturbations the phrasing of the question is intentionally altered—for instance, by introducing typographical errors (e.g., "Key Typo," "Letter Swap"), substituting words with similar meanings ("Synonym Replacement"), or rephrasing the entire sentence ("Paraphrase"). While the question varies, the "Answer Options" remain constant, consistently ranging from "1-Very important" to "4-Not important at all.". In "(10) Paraphrase x Reversal," the question is paraphrased, and simultaneously, the order of the "Answer Options" is inverted, starting with "4-Not important at all" and ending with "1-Very important."
Type Perturbation Question Answer Options Bias and Reference
Non-bias Perturbation
(5) Key Typo
nor eaca jf the following aspecto,
indicete how important it ir wn your liae.
Would bou say it is very imporcant,
rathes importano,
not very imporgant ob not impodtant at all?
Famizy
[’1=Very important ’, ’2=Rather important ’,
’3=Not very important ’, ’4=Not important at all’,
"-1=Don’t know"]
Dong et al. (2023); Moradi and Samwald (2021)
(6) Letter Swap
For each of the following sapects,
indicate how important it is in your life.
uoWld you yas it is evry important,
ratreh important, ton
very important or not important ta all? Family
Hagen et al. (2017); Moradi and Samwald (2021); Zhuang and Zuccon (2021)
(7) Keyboard Typo
For esch of rhe following aspects,
indicate how important ut is un your lide.
Would you say it ia very imporrant, rather importamt,
nit very important ir nor important ay all? Fanily
Gan et al. (2024); Zhuang and Zuccon (2021)
(8) Synonym Replacement
Crucial in life: Family For each of the following aspects,
indicate how significant it is in your life.
Would you say it is very important, rather important,
not very important or not at all important? Family
Qiang et al. (2024); Gereti et al. (2024)
(9) Paraphrase
How important is family to you?
Please rate its significance in your life
on a scale of "very important" to "not important at all".
Dong et al. (2023); Qiang et al. (2024)
Interaction
(10) Paraphrase x Reversal
How important is family to you?
Please rate its significance in your life
on a scale of "very important"
to "not important at all".
["-1=Don’t know",
’4=Not important at all’, ’3=Not very important ’,
’2=Rather important ’, ’1=Very important ’]
Dong et al. (2023)
Table 2: An exemplary perturbation scheme showing non-bias and interaction perturbations. The example is taken from the item set of category "Importance of Life Aspects". In the WVS wave 7 it is question Q1. Non-bias perturbations have variation in the question phrasing with constant answer options, while the interaction perturbation varies both.
Appendix C Results
This section summarizes the findings discussed in the main part of the work and can serve as a reference to identify other response patterns as the heatmaps contain much information regarding LLMs, perturbation type as well as additional statistical tests on the response distributions.
C.1 Robustness against Non-Bias Perturbations
This heatmap highlights to which extent the LLMs are affected by non-bias perturbations. We can see large model-specific differences.
We see that especially the smallest Llama models are responding not in a robust way. These results are consistent across the different non-bias perturbations. Especially when facing more than one perturbation in the Interaction perturbation, where both answer option scale and question phrasing were altered, the robustness drops drastically across all models .
Moreover, we identify perturbations that are less robustness-harming than others. For example, swapping letters within a word does not impact response robustness as much as typos or introducing completely different words, synonyms, or rephrasing the whole sentence.
Figure 4: Share of fully robust responses (KL divergence = 0) by model and perturbation type. Larger models (Llama-70B, Gemini) are substantially more robust than the smallest models (Llama-1B, Llama-3B).
C.2 Distance Calculation for Central Tendency and Opinion Floating Bias
This section shows how the response shifts to the central category is measured for the bias perturbations Odd/Even Scale Transformation and Priming Suffix. By calculating the differences in distances to the central point of the answer option scale we try to identify if the average distribution shifts to the central scale point. The actual differences in distance for each answer option scale type and for the two perturbations are visualized in Figure 5(a) and Figure 5(b).
To better understand how the shift towards the middle is measured, we present an anecdotal visualization in Figure 7 of the thought behind whether we observe a central tendency bias or an opinion floating bias.
In addition, Figure 6 underlines that a middle category is significantly more often chosen than assumed under a uniform, or random, distribution of responses across the scale.
Thus, a twofold analysis of not only investigating the shift of average responses between response distributions in the perturbation and original setting is important, but also the statistically assessment to make grounded claims.
By conducting a statistical binomial test, we tried to account for that (Figure 6).
(a) Models adjust their answer behavior towards the middle when the refusal category is missing (green).
(b) Models adjust their answer behavior towards the middle when a middle category is existent (green).
Figure 5: The values display the difference in mean distance of the perturbed, (a) without refusal category and (b) with middle category to the scale center. For original even scales an artificial middle category is created and vice versa to be able to compare even and odd scales with one another for every question. Thus, in an original 5-pt Likert scale the middle category is removed, whereas in a 4-pt Likert scale a middle category is added. The difference in shift of the mean response to the center is consistent across all LLMs. No changes are removed for better readability.
Figure 6: P-Values of a Binomial Test on the Middle Category in an Odd Scale. The p-values indicate a hypothesis test with a Null-Hypothesis stating that the middle category is not selected significantly more often assumed under a uniform distribution or completely random selection of answer options. However, for larger scale types, the middle category becomes much more relevant as it is significantly chosen more frequently than any other category.
Figure 7: Exemplary Difference in Distances to Scale Center of Responses to a Perturbed and Original Q&A Pair. The absolute distance is measured between the scale center and the response mean. Then, $D=D_{perturbed}-D_{original}$. A negative result indicates that the mean response in the perturbed setting is closer to the ideal scale center.
C.3 Refusal and Invalid Responses
This section should give a broader overview of the refusal rates (LLM chose the "Don’t know" answer option) and the invalid responses (e.g. a LLM did not return any valid response).
It is important to have inspect the overall return rates for all the models as these might have implications on the interpretability of the results. For example, when a model exhibits high refusal or invalid response rates, its results might be not very well interpretable as the main analysis only focused on the valid responses.
Therefore, this analysis gives insights which results are more reliable than others as for some models there are more valid responses as for others. For example, for Qwen we can see large refusal and invalid response rates, generally, but especially in sensitive thematic areas, such as Perceptions of Elections.
(a) Combined share of invalid and refusal responses by model and perturbation.
(b) Perception of Elections: Refusal and Unsuccessful Rate by Model and Perturbation.
Figure 8: Combined share of invalid and refusal responses by model and perturbation. Larger models like Llama-70B, Phi-3.5, and Mistral are highly reliable. Qwen2.5 shows high non-response rates, especially for typo-based perturbations and sensitive topics. Figure 8(b) illustrates the refusal rate in the category "Perception of Elections" separated by perturbation type and model. Large model-specific item non-responses can be identified in this thematic category indicating potential guardrails or restrictions in this domain.
C.4 Consistency of LLM Survey Responses
This section shows how consistent different LLMs respond to close-ended survey questions when facing the same Q&A pair multiple times. As explained in Section 3, we provide each model with the same Q&A pair in each perturbation stage 25 times and request a response. By running the same setting multiple times we try to identify how consistent LLMs respond generally and whether there are differences when facing the same, but syntactically incorrect (e.g. key typos, etc.), prompts.
Figure 9 highlights that the LLMs consistency in responding is not really affected by specific perturbation types. Thus, "incorrect, flawed prompts" do not increase the response inconsistency of LLMs.
However, there are again large differences between models. We see that especially the smallest Llama models and Qwen exhibit strong inconsistent responses given the same Q&A pair 25 times, whereas larger LLMs are very consistent. Nonetheless, the Llama model family seems to be more inconsistent, generally.
Figure 9: Large model-specific differences in response entropy. Little to no perturbation-specific differences. Each scale size subsumes all selected questions. This figure displays the mean entropy across all questions in that scale type for all perturbation and model combinations. Warmer colors indicate a higher average dispersion of the responses across the potential answer options. E.g., if a model answers always with the same label, the entropy is 0.
C.5 Comparison of Robustness against Perturbations by Scale Size
The following plots try to reveal in more detail the extent of robustness drop by perturbation depending on the answer option scale dimension. We identified a drop in robustness as the scale size became larger.
The responses of the smallest LLMs are generally not robust at all. However, when the scale has ten options the responses robustness plummets across all perturbations and not even a single response distribution returned in the original Q&A setting can be generated. This indicated that the smallest Llama as well as the chinese LLMs Qwen and Yi are not able to cope with perturbations, especially when facing a lot of options to choose from.
In all cases, the interaction perturbation with both question and answer option alterations leads to the largest drop in robustness across all models and scale sizes. It is striking, that larger models, especially state-of-the-art models like Gemini-1.5-Pro, cannot answer robustly given multiple perturbations.
(a) 3-point Likert Scale
(b) 4-point Likert Scale
Figure 10: Model-specific differences in fully robust responses on most perturbations on the 3 and 4-point scale. This figure shows the share of fully robust response distributions given the original response distribution and the responses based on the specific perturbation on the y-axis. Compared to 11 the robustness of responses drops when the scale size becomes larger. The smallest Llama models perform very poorly across all scales.
Researchers should take into account the scale size when generating synthetic responses from close-ended survey Q&A pairs.
(a) 5-point Likert Scale
(b) 10-point Likert Scale
Figure 11: Model-specific differences in fully robust responses on most perturbations on the 5- and 10-point scale. This figure shows the share of fully robust response distributions given the original response distribution and the responses based on the specific perturbation on the y-axis. Compared to 10 the robustness of responses drops when the scale size becomes larger. The smallest Llama models perform poorly across all scales.